Hello! can anyone please answer this problem...
Pol is 10 years younger than Greg. In 7 years he will be 10 years more than one half as old as Greg. Find their age at present.
>.Thank You.<

Answer by htmentor(1343)

Given:
Ages now: P = G - 10 [1]
Ages in 7 yrs: P + 7 = (G+7)/2 + 10 [2]
Substitute the value for P in [1] into [2]:
(G-10) + 7 = (G+7)/2 + 10
G - 3 = (G+7)/2 + 10
G/2 = 13 + 7/2
G = 33
Therefore P = 33 - 10 = 23
Ans: Pol is 23, Greg is 33